OpenStudy (anonymous):

Evaluate: 32 + (5 − 2) • 4 −

OpenStudy (nilankshi):

BODMAS

OpenStudy (anonymous):

what

OpenStudy (anonymous):

32 + 3x4 = 32 + 12 = 44